OpenStudy (anonymous):

integral e^x/sqrt(e^2x + 9)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ e ^{x} }{ e ^{2x} +9}dx\]

OpenStudy (anonymous):

looks like arctangent more or less right?

OpenStudy (anonymous):

try \(u=e^x\)

OpenStudy (anonymous):

an arc tangent? how do you know by looking?

OpenStudy (anonymous):

cause it looks like \[\frac{1}{x^2+1}\] except with a 9

OpenStudy (anonymous):

so if you put \(u=e^x\) and \(du=e^xdx\) you get \[\int \frac{du}{u^2+9}\] right away

zepdrix (zepdrix):

Make sure you remember your rules of exponents:\[\Large\sf e^{2x}=(e^x)^2\]Otherwise that denominator can look a little confusing.

OpenStudy (anonymous):

\[\LaTeX\] master

OpenStudy (anonymous):

what happens after this though?

myininaya (myininaya):

well he said to use a tan sub except you might want some constant multiple with that tan sub since you have a 9 there instead of 1